# Correspondences of the Third Kind: Camera Pose Estimation from Object Reflection

### v1

@@ Appendix A Ambiguity in The Estimation with Pixel and 3D Correspondences
that transforms surface normals of the inaccurate surface geometry to ones in the recovered normal maps in the second view. $\mathrm{G}_{2}$ is a transformation from the ground-truth geometry to the recovered normal map and $\mathrm{G}_{\hat{\eta},2}$ is one from the ground truth to the inaccurate geometry
$\mathrm{G}_{\hat{\eta},2}=\begin{pmatrix}1&0&o\\ 0&1&0\\ \mu_{\hat{\eta}}&\nu_{\hat{\eta}}&\lambda_{\hat{\eta}}\end{pmatrix}\,.$ (A.30)
@@
Similarly, we can also derive the relationship between $\hat{\eta}$ and the GBR transformation of the first view $\hat{\mathrm{G}}_{1}$. Note that, when the signs of $\eta$ and $\hat{\eta}$ are different, $\lambda_{\hat{\eta}}$ becomes negative, i.e., the surface geometry is inverted. This causes corresponding estimates of $\lambda_{1}$ and $\lambda_{2}$ to be negative thus we can exclude such a solution in practice,
Appendix B Two-Step Estimation of Relative Camera Pose
Given 3 sets of $N_{\mathrm{IM}}$ pixel correspondences, $N_{\mathrm{NM}}$ 3D correspondences, and $N_{\mathrm{RM}}$ reflection correspondences between two views, we estimate the relative rotation $\mathrm{R_{21}}$ and parameters for the generalized bas-relief transformations for both images $\mu_{1}$, $\nu_{1}$, $\lambda_{1}$, $\mu_{2}$, $\nu_{2}$, $\lambda_{2}$, under the constraints that $\lambda_{1}$ and $\lambda_{2}$ are positive. As briefly described in the main text, we achieve this by deriving a two-step algorithm. First, we obtain initial estimates of the unknown parameters using only pixel and 3D correspondences, i.e., by minimizing the sum of Eq. (14) and Eq. (17) in the main text using an off-the-shelf solver [18]. To ensure that $\lambda_{1}$ and $\lambda_{2}$ are positive, instead of directly optimizing them, we optimize their logarithms. This gives us a unique solution for $\phi$ and $\theta$, two of three Euler angles that correspond to $\mathrm{R_{21}}$, and the compound transformation matrix $\mathrm{G_{21}}\equiv\mathrm{G_{1}}^{\mathrm{-1}}\mathrm{R_{21}}\mathrm{G_{2}}^{\mathrm{T}}$. There is, however, still an ambiguity regarding $\eta$, one of the three Euler angles (Appendix A).
If $\eta$ is given in addition to $\theta$ and $\phi$, i.e., if $\mathrm{R_{21}}$ is determined, we can solve for the parameters of the generalized bas-relief transformations analytically using $\mathrm{R_{21}}$ and $\mathrm{G_{21}}$. From the definitions of $\mathrm{G_{21}}$, $\mathrm{G_{1}}$, $\mathrm{G_{2}}$, and $\mathrm{R_{21}}$, we have
$\mathrm{G_{21}}=\begin{pmatrix}\gamma_{11}&\gamma_{12}&\gamma_{13}\\ \gamma_{21}&\gamma_{22}&\gamma_{23}\\ \gamma_{31}&\gamma_{32}&\gamma_{33}\end{pmatrix}\,,$ (B.1)
$\gamma_{11}$ $=r_{11}-r_{31}\frac{\mu_{1}}{\lambda_{1}}\,,$ (B.2)
$\gamma_{12}$ $=r_{12}-r_{32}\frac{\mu_{1}}{\lambda_{1}}\,,$ (B.3)
$\gamma_{13}$ $=r_{11}\mu_{2}+r_{12}\nu_{2}+r_{13}\lambda_{2}-\mu_{1}\gamma_{33}\,,$ (B.4)
$\gamma_{21}$ $=r_{21}-r_{31}\frac{\nu_{1}}{\lambda_{1}}\,,$ (B.5)
$\gamma_{22}$ $=r_{22}-r_{32}\frac{\nu_{1}}{\lambda_{1}}\,,$ (B.6)
$\gamma_{23}$ $=r_{21}\mu_{2}+r_{22}\nu_{2}+r_{23}\lambda_{2}-\nu_{1}\gamma_{33}\,,$ (B.7)
$\gamma_{31}$ $=\frac{r_{31}}{\lambda_{1}}\,,$ (B.8)
$\gamma_{32}$ $=\frac{r_{32}}{\lambda_{1}}\,,$ (B.9)
$\gamma_{33}$ $=\frac{1}{\lambda_{1}}(r_{31}\mu_{2}+r_{32}\nu_{2}+r_{33}\lambda_{2})\,,$ (B.10)
where $r_{ij}$ is the (i,j) element of $\mathrm{R_{21}}$. From Eq. B.8 and Eq. B.9, we have
$\lambda_{1}^{2}=\frac{r_{31}^{2}+r_{32}^{2}}{\gamma_{31}^{2}+\gamma_{32}^{2}}\,.$ (B.11)
@@ Appendix B Two-Step Estimation of Relative Camera Pose
$\lambda_{1}=\sqrt{\frac{r_{31}^{2}+r_{32}^{2}}{\gamma_{31}^{2}+\gamma_{32}^{2}}}\,.$ (B.12)
@@
$\det\left(\mathrm{G_{21}}\right)=\det\left(\mathrm{G_{1}}^{\mathrm{-T}}\right)\det\left(\mathrm{R_{21}}\right)\det\left(\mathrm{G_{2}}^{\mathrm{T}}\right)=\frac{\lambda_{2}}{\lambda_{1}}\,,$ (B.13)
we also have a solution for $\lambda_{2}$
$\lambda_{2}=\det\left(\mathrm{\mathrm{G_{21}}}\right)\lambda_{1}\,.$ (B.14)
@@
As Eqs. B.2, B.3, B.4, B.5, B.6, B.7 and B.10 are linear equations for the remaining unknown parameters, we can analytically solve these equations in the least-squares sense. Note that, we empirically found that there is a solution for these equations only when the sign of the given $\eta$ and one of the ground truth are identical. This is consistent with the discussion in Appendix A.
Based on these, in the second step, we discretize $\eta$ in 180 angles, solve for the other parameters analytically, and check if the obtained parameters are consistent with the reflection correspondences.
Appendix C Implementation Details
C.1 Feature Extraction from Reflecance Maps
Similar to the feature extraction network for normal maps, we train another feature extraction network for correspondence detection from reflectance maps. We use a set of two-view synthetic reflectance maps as training data. To feed the reflectance maps into a 2D convolutional neural feature extraction network, we convert them to 2D images using angular fisheye projection. For each view, the network takes a projected reflectance map as inputs and outputs a corresponding 2D feature map. Using ground truth correspondences between surface normals $\mathbf{n_{1}^{i}}\leftrightarrow\mathbf{n_{2}^{i}}$, we impose the InfoNCE loss [17] (Eq. (20) in the main text) to ensure that features of pixels (normal orientations) that correspond to the same light source become similar.
During training, we augment the synthetic reflectance maps using Eq. (12) in the main text. Based on Eq. (4) in the main text, we also change the ground truth correspondences $\mathbf{n_{1}^{i}}\leftrightarrow\mathbf{n_{2}^{i}}$ to $\mathrm{G_{1}}^{\mathrm{-T}}\mathbf{n_{1}^{i}}\leftrightarrow\mathrm{G_{2}}^{\mathrm{-T}}\mathbf{n_{2}^{i}}$.
At inference time, similar to the detection of 3D correspondences, we detect reflection correspondences by brute-force matching and filter them with the ratio test by Lowe [10].
C.2 Network Architecture
We use a UNet architecture described in [21] (Tab. 5) to implement the feature extraction networks for correspondence detection. Each of the feature extraction network consists of a series of two UNets. We set the number of output channels of the UNets to 36.
C.3 DeepShaRM for Single-View Estimation
@@
While DeepShaRM [21] is originally designed for multi-view images (with known camera poses), we found that its joint shape and reflectance map estimation framework can work even with a single-view image. To reduce computational burden, instead of using a 3D grid of a signed distance function, we represent the surface geometry as a 2D grid of the surface depths. As the single-view estimation suffers from the bas-relief ambiguity, during the joint optimization of surface geometry and a reflectance map, we impose a regularization loss
$L_{\mathrm{reg}}=\left\{\begin{pmatrix}0\\ 0\\ l\end{pmatrix}-\frac{1}{|P|}\sum_{u,v\in P}\mathbf{N}[u,v]\right\}^{2}\,,$ (C.1)
@@
where $\mathbf{N}[u,v]$ is a normal map extracted from the depth grid, $P$ is a set of pixels inside the object, and $|P|$ is the number of such pixels. As we can assume that the surface is roughly fronto-parallel, we enforce the x and y components of the average surface normal orientation to be zero. We also encourage its z component to be identical to its average value in training data $l$. As the value of the z component corresponds to flatness of the surface (i.e., higher values means that almost all surface normals are facing the camera), this loss ensures that the surface is not extremely flat or sharp. For the rest, we follow the method described in the paper [21].
Appendix D Evaluation on Real Images
We quantitatively and qualitatively evaluate our method on real-world two-view images from the nLMVS-Real dataset [20]. We compared our results with ones of COLMAP [16] and SAMURAI [3]. We found that COLMAP completely fails on our inputs, i.e., two-view cropped images that covers only the target object. We used 11 view uncropped images that capture not only the target object but also textured object around the target (i.e., ChArUco boards) as inputs to COLMAP. Note that the uncropped images are provided by the authors of the dataset [20] as RAW data. As SAMURAI requires coarse initial camera poses, we use the same camera pose as the initial value for all input views.
Table 2 and Fig. 10 show quantitative and qualitative results. Note that COLMAP completely failed on the “Horse” object. Note also that, as SAMURAI failed to extract a 3D mesh model from their volumetric geometry representation, we instead show a normal map in the first view for SAMURAI. While existing methods fail on the extremely challenging inputs, our method successfully recovers plausible camera poses, surface geometry, and reflectance maps.
Figure 10: Reflectance maps (RMs), surface geometry, relative camera poses, and correspondences recovered from two-view real-world images.
Figure 11: Effectiveness of the use of reflection correspondences.
Planck Horse Bunny
COLMAP [16] (*) 15.7 deg N/A 24.5 deg
SAMURAI [3] 5.6 deg 24.3 deg 30.5 deg
w/o RM 3.8 deg 20.2 deg 11.9 deg
Ours 1.9 deg 12.5 deg 7.2 deg
Table 2: Camera pose estimation accuracy on real-world images in the nLMVS-Real dataset [20]. (*) COLMAP [16] uses 11 view uncropped images as inputs. Please see the text for details.
D.1 Effectiveness of Reflection Correspondences
We also compared with our method “w/o RM”, a variant of our method that only leverages pixel and 3D correspondences. To obtain a unique solution without using reflection correspondences, we ignore the bas-relief ambiguity and recovered the relative rotation based on Eq. (3) in the main text. For detection of inlier 3D correspondences, we use the same RANSAC-based algorithm as one in the first-step of our camera pose estimation method.
Table 2 and Fig. 11 show quantitative and qualitative results. Note that the joint shape and camera pose recovery without reflection correspondences completely failed on the “Horse” object. In Tab. 2, we report the accuracy of our method “w/o RM” without the joint estimation for the “Horse” object. Without reflection correspondences, the recovered camera poses can be wrong and the recovered geometry can be distorted due to the bas-relief ambiguity.

### v2

@@ Appendix 0.C Ambiguity in The Estimation with Conventional Correspondences
that transforms surface normals of the inaccurate surface geometry to those of the recovered normal maps in the second view. $\mathrm{G}_{2}$ is the transformation from the ground-truth geometry to the recovered normal map and $\mathrm{G}_{\hat{\eta},2}$ is that from the ground truth to the inaccurate geometry
$\mathrm{G}_{\hat{\eta},2}=\begin{pmatrix}1&0&o\\ 0&1&0\\ \mu_{\hat{\eta}}&\nu_{\hat{\eta}}&\lambda_{\hat{\eta}}\end{pmatrix}\,.$ (0.C.30)
@@
Similarly, we can also derive the relationship between $\hat{\eta}$ and the GBR transformation of the first view $\hat{\mathrm{G}}_{1}$. Note that, when the signs of $\eta$ and $\hat{\eta}$ are different, $\lambda_{\hat{\eta}}$ becomes negative, i.e., the surface geometry is inverted. This causes corresponding estimates of $\lambda_{1}$ and $\lambda_{2}$ to be negative. We can easily exclude such a solution in practice.
Appendix 0.D Analytical Decomposition of $\mathrm{G}_{21}$
As we discussed in the main text and in Appendix 0.C, given just the pixel and 3D correspondences, we can only obtain unique solutions for the combined transformation $\mathrm{G}_{21}\equiv\mathrm{G_{1}}^{-\mathrm{T}}\mathrm{R_{21}}\mathrm{G_{2}}^{\mathrm{T}}$ and the two Euler angles $(\theta,\phi)$. However, if the remaining Euler angle $\eta$ is given in addition to $\theta$ and $\phi$, i.e., if $\mathrm{R_{21}}$ is determined, we can solve for the parameters of the generalized bas-relief transformations analytically. For this, we derive a two-step relative rotation estimation algorithm as detailed in Appendix 0.E.
From the definitions of $\mathrm{G_{21}}$, $\mathrm{G_{1}}$, $\mathrm{G_{2}}$, and $\mathrm{R_{21}}$, we have
$\mathrm{G_{21}}=\begin{pmatrix}\gamma_{11}&\gamma_{12}&\gamma_{13}\\ \gamma_{21}&\gamma_{22}&\gamma_{23}\\ \gamma_{31}&\gamma_{32}&\gamma_{33}\end{pmatrix}\,,$ (0.D.1)
$\gamma_{11}$ $=r_{11}-r_{31}\frac{\mu_{1}}{\lambda_{1}}\,,$ (0.D.2)
$\gamma_{12}$ $=r_{12}-r_{32}\frac{\mu_{1}}{\lambda_{1}}\,,$ (0.D.3)
$\gamma_{13}$ $=r_{11}\mu_{2}+r_{12}\nu_{2}+r_{13}\lambda_{2}-\mu_{1}\gamma_{33}\,,$ (0.D.4)
$\gamma_{21}$ $=r_{21}-r_{31}\frac{\nu_{1}}{\lambda_{1}}\,,$ (0.D.5)
$\gamma_{22}$ $=r_{22}-r_{32}\frac{\nu_{1}}{\lambda_{1}}\,,$ (0.D.6)
$\gamma_{23}$ $=r_{21}\mu_{2}+r_{22}\nu_{2}+r_{23}\lambda_{2}-\nu_{1}\gamma_{33}\,,$ (0.D.7)
$\gamma_{31}$ $=\frac{r_{31}}{\lambda_{1}}\,,$ (0.D.8)
$\gamma_{32}$ $=\frac{r_{32}}{\lambda_{1}}\,,$ (0.D.9)
$\gamma_{33}$ $=\frac{1}{\lambda_{1}}(r_{31}\mu_{2}+r_{32}\nu_{2}+r_{33}\lambda_{2})\,,$ (0.D.10)
where $r_{ij}$ is the $(i,j)$ element of $\mathrm{R_{21}}$. From Eq. 0.D.8 and Eq. 0.D.9, we have
$\lambda_{1}^{2}=\frac{r_{31}^{2}+r_{32}^{2}}{\gamma_{31}^{2}+\gamma_{32}^{2}}\,.$ (0.D.11)
@@ Appendix 0.D Analytical Decomposition of $\mathrm{G}_{21}$
$\lambda_{1}=\sqrt{\frac{r_{31}^{2}+r_{32}^{2}}{\gamma_{31}^{2}+\gamma_{32}^{2}}}\,.$ (0.D.12)
@@
$\det\left(\mathrm{G_{21}}\right)=\det\left(\mathrm{G_{1}}^{\mathrm{-T}}\right)\det\left(\mathrm{R_{21}}\right)\det\left(\mathrm{G_{2}}^{\mathrm{T}}\right)=\frac{\lambda_{2}}{\lambda_{1}}\,,$ (0.D.13)
we also have a solution for $\lambda_{2}$:
$\lambda_{2}=\det\left(\mathrm{\mathrm{G_{21}}}\right)\lambda_{1}\,.$ (0.D.14)
@@
As Eqs. 0.D.2, 0.D.3, 0.D.4, 0.D.5, 0.D.6, 0.D.7 and 0.D.10 are linear equations for the remaining unknown parameters, we can analytically solve these equations in the least-squares sense. Note that, we empirically found that there is a solution for these equations only when the sign of the given $\eta$ and one of the ground truth are identical. This is consistent with the discussion in Appendix 0.C.
Appendix 0.E Two-Step Relative Rotation Estimation
1: // Estimation of $\mathrm{G}_{21}$ and $(\phi,\theta)$ 2: // with pixel and 3D correspondences 3:$bestScore1\leftarrow 0$ 4:for $i\leftarrow 1$ to $L_{1}$ do 5:    $S_{1}\leftarrow$ random $M_{1}$ indices from $N_{\mathcal{N}}$ indices 6:    $\hat{\mathrm{G}}_{21},\hat{\phi},\hat{\theta}\leftarrow\operatorname*{arg\,min}_{\mathrm{G}_{21}}f_{1}(S_{1})$ 7:    if $g_{1}\left(\hat{\mathrm{G}}_{21},\hat{\phi},\hat{\theta}\right)>bestScore1$  then 8:       $bestScore1\leftarrow g_{1}\left(\hat{\mathrm{G}}_{21},\hat{\phi},\hat{\theta}\right)$ 9:       $bestModel1\leftarrow\hat{\mathrm{G}}_{21},\hat{\phi},\hat{\theta}$ 10:    end if 11:end for 12:$\hat{\mathrm{G}}_{21},\hat{\phi},\hat{\theta}\leftarrow bestModel1$ 13: // Decomposition of $\mathrm{G}_{21}$ 14: // with $g_{2}(\cdot)$ which exploits reflection correspondences 15:$bestScore2\leftarrow 0$ 16:for $\eta\leftarrow$ 1 deg to 359 deg do 17:    $\hat{\mathrm{R}}_{21},\hat{\mathrm{G}}_{1},\hat{\mathrm{G}}_{2}\leftarrow$ decomposition($\hat{\mathrm{G}}_{21};\hat{\phi},\eta,\hat{\theta}$) 18:    if $g_{2}\left(\hat{\mathrm{R}}_{21},\hat{\mathrm{G}}_{1},\hat{\mathrm{G}}_{2}\right)>bestScore2$ then 19:       $bestScore2\leftarrow g_{2}\left(\hat{\mathrm{R}}_{21},\hat{\mathrm{G}}_{1},\hat{\mathrm{G}}_{2}\right)$ 20:       $bestModel2\leftarrow\hat{\mathrm{R}}_{21},\hat{\mathrm{G}}_{1},\hat{\mathrm{G}}_{2}$ 21:    end if 22:end for 23:return  $bestModel2$
Algorithm 1 Relative Rotation Estimation
Algorithm 1 shows a pseudo code of the two-step relative rotation estimation from the three types of correspondences. In the first step, using $N_{\mathcal{N}}$ pixel and $N_{\mathcal{N}}$ 3D correspondences, we obtain estimates of the combined transformation $\mathrm{G}_{21}$ and the two Euler angles $(\phi,\theta)$ based on RANSAC [9]. For each of the $L_{1}$ iterations, we first build a set of randomly sampled $M_{1}$ indices for the pixel and 3D correspondences. We then obtain estimates of the combined transformation $\hat{\mathrm{G}}_{21}$ and the Euler angles $(\hat{\phi},\hat{\theta})$ by minimizing an objective function
$f_{1}(S_{1})=f_{\mathcal{I}}(S_{1})+f_{\mathcal{N}}^{(12)}(S_{1})+f_{\mathcal{N}}^{(21)}(S_{1})\,,$ (0.E.1)
with respect to the relative rotation and parameters of the generalized bas-relief transformations ($\mu_{1},\nu_{1},\lambda_{1},\mu_{2},\nu_{2},\lambda_{2}$). $f_{\mathcal{I}}(S_{1})$ and $f_{\mathcal{N}}^{(jk)}(S_{1})$ are similar to those in Eq. (13) in the main text. We compute them using only the sampled correspondences
$f_{\mathcal{I}}(S_{1})=\frac{1}{{N_{\mathcal{N}}}}\sum_{i\in S_{1}}\left(t_{\phi}^{i}-t_{\theta}^{i}\right)^{2}\,,$ (0.E.2)
$f_{\mathcal{N}}^{(jk)}(S_{1})=\frac{1}{{N_{\mathcal{N}}}}\sum_{i\in S_{1}}\left\|{\bf N}_{j}[u_{j}^{i},v_{j}^{i}]-{\text{Norm}}(\mathrm{G}_{kj}{\bf N}_{k}[u_{k}^{i},v_{k}^{i}])\right\|^{2}\,,$ (0.E.3)
with
$t_{\phi}^{i}=u_{1}^{i}\cos\phi+v_{1}^{i}\sin\phi\,,$ (0.E.4)
$t_{\theta}^{i}=u_{2}^{i}\cos\theta+v_{2}^{(i)}\sin\theta\,.$ (0.E.5)
In practice, we achieve this minimization by using an off-the-shelf solver for nonlinear optimization [31, 5]. We evaluate the goodness of the estimates using a function
$g_{1}\left({\mathrm{G}}_{21},\phi,\theta\right)=\frac{1}{N_{\mathcal{N}}}\sum_{i=1}^{N_{\mathcal{N}}}\exp\left(-\frac{p_{i}^{2}}{T_{\mathcal{I}}^{2}}\right)\exp\left(-\frac{\left(q_{i}^{(12)}\right)^{2}+\left(q_{i}^{(21)}\right)^{2}}{T_{\mathcal{N}}^{2}}\right)\,,$ (0.E.6)
where $p_{i}$ and $q_{i}^{(jk)}$ are residuals with respect to the pixel and the 3D correspondences
$p_{i}=t_{\phi}^{i}-t_{\theta}^{i}\,,$ (0.E.7)
$q_{i}^{(jk)}=\arccos\left({\bf N}_{j}[u_{j}^{i},v_{j}^{i}]\cdot{\text{Norm}}(\mathrm{G}_{kj}{\bf N}_{k}[u_{k}^{i},v_{k}^{i}])\right)\,,$ (0.E.8)
and $T_{\mathcal{I}}$ and $T_{\mathcal{N}}$ are corresponding thresholds. This function computes the approximated number of correspondences that are consistent with the obtained estimates. We select a set of the estimates $(\hat{\mathrm{G}}_{21},\hat{\phi},\hat{\theta})$ that maximizes this function.
In the second step, we discretize $\eta$ in 360 angles and, for each $\eta$, decompose the estimate of the combined transformation $\hat{\mathrm{G}}_{21}$ into the relative rotation $\mathrm{R}_{21}$ and the GBR transformations $\hat{\mathrm{G}}_{1}$ and $\hat{\mathrm{G}}_{2}$ by using the method in Appendix 0.D. We then select a set of the estimates that maximizes
$g_{2}\left(\mathrm{R}_{21},\mathrm{G}_{1},\mathrm{G}_{2}\right)=s_{\mathrm{reg}}\frac{1}{N_{\mathcal{R}}}\sum_{i=1}^{N_{\mathcal{R}}}\exp\left(-\frac{\left(r_{i}^{(12)}\right)^{2}+\left(r_{i}^{(21)}\right)^{2}}{T_{\mathcal{R}}^{2}}\right)\,,$ (0.E.9)
where $r_{i}^{(jk)}$ is the residual with respect to the reflection correspondences
$r_{i}^{(jk)}=\arccos\left({\bf N}_{j}[u_{j}^{i},v_{j}^{i}]\cdot\Omega^{(jk)}({\bf N}_{k}[u_{k}^{i},v_{k}^{i}])\right)\,,$ (0.E.10)
and $T_{\mathcal{R}}$ is a corresponding threshold.
In practice, we set $L_{1}$, $T_{\mathcal{I}}$, $T_{\mathcal{N}}$, and $T_{\mathcal{R}}$ to be 200, 5 pixels, 20 degrees, and 10 degrees, respectively. We set $M_{1}$ to be 4 according to empirical evaluation below.
0.E.1 Empirical Evaluation about The Number of Correspondences
As mentioned in Section 3.4 in the main text, we conduct empirical evaluation of the number of correspondences required for each step of the two-step algorithm. We created 10000 sets of synthetic correspondences by randomly sampling surface normal orientations, relative rotation matrices, and parameters of the GBR transformation. For each set of correspondences, we optimized all the unknown parameters with all available correspondences (i.e., Eq. (13) in the main text) using an off-the-shelf solver [31, 5] and checked if the estimates are accurate.
Figure E.17: Empirical analysis of the number of correspondences required for optimization of the camera pose and the two GBR transforms of two images. See text for details.
The blue line in Fig. E.17 shows the relationship between the number of correspondences in normal maps (pixel and 3D correspondences) and rate of failure cases where the error (Frobenius norm) between estimated combined transformation $\mathrm{G_{21}}$ and the ground truth is higher than 0.01. The results show that four correspondences are required for the estimation of $\mathrm{G_{21}}$. The orange line in Fig. E.17 shows the relationship between the number of reflection correspondences and rate of failure cases where the error between the estimated camera pose and the ground truth is higher than 0.1 degree. In this experiment, we used four pixel and four 3D correspondences in addition to the reflection correspondences. One reflection correspondence resolves the remaining ambiguity. Note that, the leftmost point in the orange line shows that, without reflection correspondences, we cannot decompose the combined transformation $\mathrm{G_{21}}$. We hope to derive a theoretical justification of these results in our future work.
Appendix 0.F Single-View Estimation with DeepShaRM [36]
As explained in the main paper, given posed multi-view images and an initial coarse estimate of surface geometry (e.g., a visual hull or a bounding sphere) represented by a 3D grid of a signed distance function, DeepShaRM [36] can jointly recover geometry, reflectance maps, and normal maps by alternating between (1) learning-based reflectance map estimation for each view from an image and the geometry estimate, (2) learning-based surface normal estimation for each view from an image and the estimated reflectance map, and (3) multi-view geometry optimization with the estimated surface normals. Note that the learning-based reflectance map estimation and the learning-based surface normal estimation are trained using synthetic images in a supervised manner.
@@ Appendix 0.F Single-View Estimation with DeepShaRM [36]
While DeepShaRM [36] is originally designed for posed multi-view images, we found that its joint shape and reflectance map estimation framework can work even with a single-view image. To reduce the computational cost, instead of using a 3D grid of a signed distance function, we represent the surface geometry as a 2D grid of the surface depths. As the single-view estimate suffers from the bas-relief ambiguity, during the geometry optimization with the estimated surface normals, we impose a regularization loss
$L_{\mathrm{reg}}=\left\{\begin{pmatrix}0\\ 0\\ l\end{pmatrix}-\frac{1}{|P|}\sum_{u,v\in P}\mathbf{N}[u,v]\right\}^{2}\,,$ (0.F.1)
@@
where $\mathbf{N}[u,v]$ is a normal map extracted from the depth grid, $P$ is a set of pixels inside the object, and $|P|$ is the number of such pixels. As we can assume that the surface is roughly fronto-parallel, we enforce the x and y components of the average surface normal orientation to be zero. We also encourage its z component to be identical to its average value in training data $l$. As the value of the z component corresponds to the flatness of the surface (i.e., higher values correspond to almost all surface normals facing the camera), this loss ensures that the surface is not extremely flat or sharp. For the remaining estimation, we follow the method described in [36].
Appendix 0.G Learning-Based Correspondence Detection
Figure G.18: We train a deep neural network to extract view-independent, per-pixel features for correspondence matching. Contrastive learning using InfoNCE loss [25] with data augmentation with the GBR transform enables feature extraction for distortion-robust correspondence matching.
Figure G.18 depicts how we train a feature extraction network for correspondence detection from normal maps. As explained in Section 3.5 in the main text, using synthetic normal maps and corresponding ground truth matches as training data, we train the network by contrastive learning. We also augment the input normal maps by generalized bas-relief (GBR) transformations so that the network can learn to extract features robust to the estimation errors caused by the bas-relief ambiguity.
0.G.1 Network Architecture
We use the UNet-based architecture described in [36] (Tab. 5) to implement the feature extraction networks for correspondence detection. Each of the feature extraction networks consists of a series of two UNets. We set the number of output channels of the UNets to 36.
Appendix 0.H Estimation for Scenes with Textured and Textureless Objects
Figure H.19: Overview of relative camera pose recovery from a sparse set of images capturing scenes consisting both textured and textureless, non-Lambertian objects. The obtained relative camera pose (“Refined Relative Pose”) is further improved by alternating between the multi-view DeepShaRM and our pose estimation method.
Figure H.19 depicts the overall framework for relative camera pose estimation from a sparse set of images capturing scenes consisting both textured and textureless, non-Lambertian objects. We first detect pixel correspondences directly from images using an existing method [17] and recover the relative camera pose from them by using a conventional structure-from-motion method for a perspective model. We then recover reflectance maps , normal maps, and surface geoemtry from the (initial) estimate of the relative camera pose and cropped images that capture the textureless, non-Lambertian object using DeepShaRM [36]. We detect reflection correspondences on the recovered reflectance maps and leverage them to obtain a refined estimate of the relative rotation and a corresponding relative translation vector.
In this setting, to avoid relying on conventional correspondences on the textureless, non-Lambertian surfaces, we do not detect 3D correspondences and do not recover the combined transformation $\mathrm{G}_{21}$. We can, however, still apply the second step of our relative rotation estimation algorithm (Appendix 0.E) by assuming that $\mathrm{G}_{21}$ is identical to the relative rotation $\mathrm{R}_{21}$ estimated by the SfM method. This is reasonable because multi-view consolidation of DeepShaRM ensures that discrepancies between surface normals estimated for the first view and those for the second view (i.e., the combined transformation) can be approximated by the input relative rotation.
This enables us to apply our method to textureless, non-Lambertian objects whose pixel and 3D correspondences are ambiguous. Note that, even in this setting, if we could detect 3D correspondences on the textureless, non-Lambertian surface, they can be beneficial especially for avoiding solutions that invert the surface geometry. We hope to exploit pixel and 3D correspondences detected from different objects simultaneously in our future work.
Similar to Section 3.6 in the main text, we further improve the camera pose estimate by alternating between the multi-view reflectance map recovery by DeepShaRM and our camera pose estimation method.
0.H.0.1 Implementation of Structure-from-Motion
We found that off-the-shelf structure-from-motion methods for perspective projection (e.g., an implementation in [4]) often completely fail when cameras are almost orthographic. For a fair comparison, we improved the accuracy of the SfM baseline method by leveraging the orthographic assumption. That is, we first estimate the two Euler angles $(\phi,\theta)$ based on the orthographic model. We find a solution that maximizes the number of inlier correspondences whose reprojection errors (Eq. 0.E.7) are lower than 5 pixels. We then estimate the other camera parameters by minimizing reprojection errors based on the perspective model.
Appendix 0.I Dataset Details
Figure I.20: Example synthetic training images from the training set of the nLMVS-Synth dataset [35]
Figure I.20 shows example synthetic training images from the training set of the nLMVS-Synth dataset [35] which we use for training the deep networks in our method. As explained in the main text, The training shapes are composed of primitive shapes (ellipsoids, cubes, cylinders) augmented with random height fields [34]. 94 materials and 2685 environmental maps from existing databases [21, 10, 38] are used for rendering.
(a) Illuminations
(b) BRDFs
(c) Shapes
Figure I.21: Synthetic test images in the nLMVS-Synth dataset [35] and some original ones are rendered using unseen illumination maps [8, 14], unseen BRDFs [21], and unseen shapes [15, 28, 7].
Figure I.21 shows unseen illumination maps, unseen BRDFs, and unseen shapes used for rendering the synthetic test images in the nLMVS-Synth dataset [35] and some rendered by ourselves. We excluded images in the nLMVS-Synth that are rendered from less specular BRDFs (e.g., “Rubber”) or monotonous illumination maps (e.g., “Uffizi”) as appearance caused by them are almost independent of surface normals and extremely challenging to both our method and existing methods. Also, we could use only images of 2 shapes (“Bunny” and “Planck”) mainly because conventional correspondences on the other shapes (e.g., “Sphere”) are ambiguous. Instead, we rendered images of a cow model [7] by ourselves.
Figure I.22: We use an object with ChArUco patterns for calibration of real images captured by ourselves.
Figure I.22 shows how we calibrate the ground truth camera poses for real images captured by ourselves. We used an object with ChArUco patterns to detect 3D-2D point correspondences, and calibrated the camera parameters by using an off-the-shelf solver [4] for the Perspective-n-Point (PnP) pose computation problem.